OpenStudy (anonymous):

What is the approximate area of a circle with a radius of 14 centimeters? Use π = 3.14. A. 2461.8 square centimeters B. 615.4 square centimeters C. 153.9 square centimeters D. 87.9 square centimeters Thanks! ☺

OpenStudy (shadowlegendx):

Lila do you know the formula to solve for the area of a circle?

OpenStudy (eyust707):

Use the same formula I taught you \[Area = \pi * r * r\]

OpenStudy (mrnood):

I guess you have in oyur study a formula for th earea of a circle It has only pi and r in the formula. Using r = 14 and pi = 3.14 and a calculator and the formula - you can work it out....

OpenStudy (eyust707):

You got the last one I have a feeling you can get this on too!

OpenStudy (shadowlegendx):

\[π\times 14^{2}\]

OpenStudy (shadowlegendx):

\[3.14 \times 14^{2}\]

OpenStudy (mrnood):

@ShadowLegendX Maybe the rest should be done by lila????

OpenStudy (shadowlegendx):

Follow orders of operations (PEMDAS) and solve for \[14^{2}\] first

OpenStudy (shadowlegendx):

A person can not do a problem that they have never seen done before

OpenStudy (mrnood):

given ALL the information above they should make an effort

OpenStudy (shadowlegendx):

They atleast need assistance with the first couple steps

OpenStudy (anonymous):

Thanks @ShadowLegendX @eyust707 I remember the formula. So I do the same with this one?

OpenStudy (anonymous):

@MrNood Sorry my internet was acting up, that's why I wasn't making an "effort"

OpenStudy (mrnood):

np - but you have all you need now to calculate this

OpenStudy (shadowlegendx):

Lila what did you get?

OpenStudy (anonymous):

Okay what if i was still confused and couldn't solve it

OpenStudy (anonymous):

Still calculating TI-32 is acting up now

OpenStudy (mrnood):

@ShadowLegendX The real world is FULL of problems that we've never seen before. The purpose of study is to learn how to extrapolate what we DO know to interpret that which we don't know

OpenStudy (shadowlegendx):

Lila you just need to solve for \[3.14 \times 14^{2}\]

OpenStudy (shadowlegendx):

3.14 is Pi(π) \[14^{2}\] is the radius

OpenStudy (mrnood):

I suggest you do it in your head - it says ESTIMATE You can do 14 x 14 in your head - then multiply by 3 - see which answer is closest.

OpenStudy (anonymous):

@ShadowLegendX I got 615.44

OpenStudy (anonymous):

Thanks Mr. Nood

OpenStudy (shadowlegendx):

Thats correct :)

OpenStudy (anonymous):

Thanks

OpenStudy (shadowlegendx):

Np :)